# Edge of Many-Body Quantum Chaos in Quantum Reservoir Computing

### v1

@@ Abstract
The most pronounced trend is observed for $\Delta t_{\mathrm{in}}=50$ [Fig. 4(a)]. When $\kappa_{2}/J_{4}$ is varied, the NARMA performance exhibits a clear peak at a coupling strength just before the onset of the many-body quantum chaotic regime. This peak provides another manifestation of the “edge of many-body quantum chaos”, in this case in the parameter domain. By contrast, for the short input interval $\Delta t_{\mathrm{in}}=1$, varying the coupling strength yields no performance enhancement [Fig. 4(b)]. At such a short time-scale, the QRC does not reflect chaotic characteristics, so the performance changes only gradually, without any distinct peak at the chaos boundary.
We remark that the specifics of performance enhancement are task-dependent. For instance, in the higher-order tasks detailed in the Supplementary Material, both the position and strength of performance peaks vary with the order [86]. Nevertheless, the fundamental trend remains robust: the most pronounced improvements appear at the edge of many-body quantum chaos. Moreover, since the NARMA task captures nonlinear temporal dynamics that mirror those in real-world systems, the performance boosts shown on this task strongly suggest that similar improvements will extend to a broad class of practical temporal processing applications.
@@
Discussions. —Recent advances in QRC have yielded related performance enhancements, and we believe that a concise comparison highlights the significance of our contributions. One of the authors and his collaborators identified a performance peak near the operational regime boundary in their feedback-driven QRC protocol [32]; however, this enhancement appears predominantly due to classical instabilities in feedback connections, rather than intrinsically quantum phenomena. In the purely quantum regime, Martínez-Peña et al. demonstrated enhanced performance near the ergodic-localization transition in the spin-based QRC [20, 23, 33]. While this finding highlights the importance of dynamical phases in QRC, the role of many-body quantum chaos, and its relation to RMT, remain largely unexplored. Recently, Llodrà et al. analyzed QRC performance in connection with quantum chaos at the quantum phase transition between superfluid and Mott insulator in a clean system, which is complementary to our study on random systems [34]. From a standpoint of temporal domain, investigations into quantum chaos remain sparse: Vetrano et al. explored the role of scrambling times, though within a different framework (quantum extreme learning machines) that does not consider memory of sequential inputs [88].
@@
Building upon these studies, our work clarifies the key characteristics that endow QRC with high computational performance. We have extended the concept of the “edge of chaos” into the quantum realm, investigating how many-body quantum chaos influences QRC in both temporal and parametric domains. The SYK model—a canonical many-body quantum chaotic system—serves as an ideal testbed. Through a unified analysis, we have revealed two distinct optimal regimes: one is the “temporal edge”, which emerges before the Thouless time, and the other is the “parametric edge”, which appears near the quantum chaotic-to-integrable transition. These findings integrate existing knowledge and establish a general principle: optimal QRC performance occurs at the “edge of many-body quantum chaos”. Our study therefore provides clear design principles for engineering quantum reservoirs, unlocking advanced information processing with quantum many-body systems.
@@ I.1 Level statistics
As illustrated in Fig. 2(b) of the main text, continuously interpolating between the SYK4 and SYK2 models drives the system from a many-body quantum chaotic phase into an integrable regime. To elucidate this behavior in greater detail, Fig. S2(a) presents the full distribution $P(r)$ for several intermediate values of the coupling ratio $\kappa_{2}/J_{4}$, extending the results presented in Fig. 2(a) of the main text. As $\kappa_{2}/J_{4}$ increases, $P(r)$ evolves smoothly from the WD (GUE) statistics toward the Poisson distribution. In Fig. S2(b), we plot the mean ratio $\langle r\rangle$ as a function of $\kappa_{2}/J_{4}$ for various system sizes at $N_{p}=\lfloor N/2\rfloor$. The chaotic-integrable transition is consistently observed, although a precise determination of the critical coupling $\kappa_{2}^{c}$ would require even larger system sizes [81]. For the purposes of this work, we provisionally define the boundary of the quantum chaotic regime by the coupling strength $\kappa_{2}^{b}$ at which $\langle r\rangle$ first deviates from its RMT value. This reference point is used to discuss the edge of many-body quantum chaos in Figs. 4(a) and 4(b) of the main text. We note that $\kappa_{2}^{b}$ and the true critical point $\kappa_{2}^{c}$ are expected to coincide in the thermodynamic limit, removing any ambiguity in the position of the parametric edge of many-body quantum chaos.
@@ I.2 Spectral form factor
Figure S3: (a), (b) The SFF for the SYK4 model and the SYK2 model, shown for various system sizes. Data are averaged over $20000$ realizations for $N=6,7,8$, $10000$ for $N=9$, and $2500$ for $N=10$. (c) The plateau time $t_{\mathrm{p}}$ of the SFF as a function of the system size $N$. The blue and red dashed lines indicate reference scaling behaviors, $\mathcal{N}=\binom{N}{N_{p}}$ and $N$, respectively.
@@
where the sum runs over the eigenvalues with the particle number $N_{p}$, and $\mathcal{N}=\binom{N}{N_{p}}$ denotes the dimension of the corresponding Hilbert space. As discussed in the main text, the SFF exhibits universal characteristics in the quantum chaotic systems: at early times, it decreases due to the phase interference, while at sufficiently late times, it shows a linear ramp before the saturation to a plateau value. In Figs. S3(a) and S3(b), we present the SFF for the SYK4 and SYK2 models, respectively, for various system sizes. The ramp-plateau behavior is clearly demonstrated especially for the former system, confirming its many-body quantum chaotic nature. In contrast, the SFF for the latter model shows the dip and plateau over a much narrower time window. Indeed, Fig. S3(c) shows that the plateau time $t_{\mathrm{p}}$ scales with the system size as $t_{\mathrm{p}}\sim\mathcal{N}$ for the SYK4 system and as $t_{\mathrm{p}}\sim N$ for the SYK2 system [70]. This further highlights the difference between the quantum chaotic and integrable models.
@@
Figure S4: The SFF for the rescaled Hamiltonian $\mathcal{H}\rightarrow\mathcal{H}/||\mathcal{H}||$ at several coupling strength $\kappa_{2}$, with a fixed system size of $N=8$ and $J_{4}=1$. The dashed lines signify $t=1$ and $50$, in relation to the input interval $\Delta t_{\mathrm{in}}$ used in to evaluate QRC performance. The results are averaged over $20000$ realizations.
@@
When interpolating between the two models, the overall energy scale varies with the coupling strength. To compare QRC performance on an equal footing, we normalize the Hamiltonian by its spectral norm, $\mathcal{H}\rightarrow\mathcal{H}/||\mathcal{H}||$. This normalization leaves the level spacing ratio invariant, as well the chaotic nature of the system. However, it introduces a uniform prefactor in the temporal domain, affecting the SFF [Eq. (S3)] and the $\Delta t_{\mathrm{in}}$ dependency of QRC performance. Figure S4 displays the SFF for the interpolated model after the rescaling. In the quantum chaotic regime ($\kappa_{2}=0.1$, $1$), the typical ramp-plateau profiles emerge, while in the integrable regime ($\kappa_{2}=20$, $100$), the SFF dips and plateaus at much earlier times. Apart from the overall time-scale shift introduced by the Hamiltonian rescaling, these behaviors mirror those of the SYK4 and SYK2 models in Fig. 2(b) of the main text.
@@
In our QRC performance benchmarks (Fig. 4 of the main text), we selected the input time intervals $\Delta t_{\mathrm{in}}=50$ and $\Delta t_{\mathrm{in}}=1$; the corresponding times are marked by the dashed lines in Fig. S4. For the quantum chaotic regime, the longer $\Delta t_{\mathrm{in}}$ falls within the RMT ramping window, while the shorter one precedes the dip. Similarly for the integrable models, these two times effectively sandwich the dip. These observations validate both the Hamiltonian normalization procedure and our choice of $\Delta t_{\mathrm{in}}$ for probing distinct temporal regimes relevant to QRC. It is worth noting that precisely characterizing chaos in the interpolated model via the SFF typically requires a more sophisticated procedure called unfolding [82]. Unfolding recasts the energy spectrum such that it will exhibit a constant level spacing; it is not an operation directly applied to the Hamiltonian itself. We do not employ this procedure because our primary motivation is to characterize QRC performance by focusing on the properties of the original, physically implemented Hamiltonian.
@@ II Dynamics of quantum reservoir system
Figures S5(a) and S5(b) illustrate the dynamics of read-out signals for the SYK4 and SYK2 models, respectively, for input intervals of $\Delta t_{\mathrm{in}}=1$ and $\Delta t_{\mathrm{in}}=10$. The dashed lines indicate when new inputs are provided. Markers show $\langle c_{i}^{\dagger}c_{i}\rangle$ at times $v\Delta t_{\mathrm{in}}/V$ ($0\leq v\leq V$), with the filled marker specifically denoting site $1$, i.e., the input site. The dynamics at the input site change abruptly after an input, showing a larger variation range than other sites; hence, its vertical axis is displayed separately.
With a short input interval ($\Delta t_{\mathrm{in}}=1$), each new input arrives before the previous one has fully influenced the system. Thus, the resulting dynamics is relatively simple, as shown in the upper panels of both Figs. S5(a) and S5(b). In contrast, system-dependent characteristics emerge when the input interval is longer ($\Delta t_{\mathrm{in}}=10$). In the SYK4 model [lower panel of Fig. S5(a)], all readout signals tend toward quasi-stationary values, whereas in the SYK2 model [lower panel of Fig. S5(b)], the signals exhibit continuous oscillations. Notably, although both dynamics appear sufficiently rich, their QRC performance differs significantly, as demonstrated in Figs. 3(c) and 3(d) of the main text. Because the output is computed from these readout signals, this performance gap must originate from the difference in dynamics; however, the underlying reason is not immediately obvious from the raw signals in Fig. S5.
These $\Delta t_{\mathrm{in}}$-dependent behaviors raise interest in the relationship between QRC performance and scrambling time—the timescale over which information delocalizes [73]. Since scrambling relates to the propagation of encoded input information, its connection to memory performance is naturally anticipated [35]. Accordingly, the observed $\Delta t_{\mathrm{in}}$ dependence of STM performance [upper panels of Figs. 3(c) and 3(d) of the main text] could plausibly reflect the underlying scrambling timescale. Indeed, their similar scaling complicates the precise distinction between the individual contributions of scrambling time and Thouless time. However, the pronounced enhancement in nonlinear processing performance [lower panel of Fig. 3(c) of the main text] cannot be explained by scrambling alone, which is primarily associated with memory. Moreover, the convergence of QRC performance toward that of the Haar QRC model when $\Delta t_{\mathrm{in}}$ exceeds the Thouless time underscores a fundamental connection to RMT. These observations collectively suggest that the Thouless time plays the more decisive role in determining overall QRC performance. Nevertheless, the detailed relationship between scrambling time, Thouless time, and the temporal edge of many-body quantum chaos warrants further investigation.
@@ III Convergence property
Figure S6: Evolution of the distances $||\rho_{A}^{(k)}-\rho_{B}^{(k)}||$ as a function of the step $k$ for (a) the SYK4 model and (b) the SYK2 model. The left panels show the median distance (semilog scale), and the right panels show the mean distance [semilog for (a) and log-log for (b)], calculated from $500$ independent pairs of initial states $\rho_{A}^{(0)}$ and $\rho_{B}^{(0)}$. Colors indicate the input interval $\Delta t_{\mathrm{in}}$.
We here examine echo state property (ESP), a fundamental characteristic of RC systems. ESP dictates that the reservoir state should asymptotically converge when driven by the same input sequence, irrespective of its initial conditions. In other words, after sufficiently many input injections, any dependence on the initial state should vanish. In QRC paradigms, ESP is typically assessed by measuring the distance between two quantum states, $\rho_{A}^{(k)}$ and $\rho_{B}^{(k)}$. These states evolve from different initial states, $\rho_{A}^{(k=0)}$ and $\rho_{B}^{(k=0)}$, but under an identical input sequence. ESP is considered satisfied when this distance approaches 0. Under purely unitary dynamics, the distance between these states, evaluated using the Frobenius norm, remains constant over time: $||\rho_{A}^{(k)}(t)-\rho_{B}^{(k)}(t)||=||\rho_{A}^{(k)}(t^{\prime})-\rho_{B}^{(k)}(t^{\prime})||$. Thus, the specific time $t$ within a unitary evolution step is irrelevant. However, the distance does change upon the nonunitary input injection process, which involves the partial trace over site $1$ and the reinitialization with the input state $\rho_{\mathrm{in}}$. Consequently, to evaluate ESP, we monitor the state distance as a function of the input step $k$. As the system traces out the previous state and injects a new, identical input state for both trajectories, the distance between $\rho_{A}^{(k)}$ and $\rho_{B}^{(k)}$ is expected to decrease with $k$. The precise convergence behavior depends on the state before the input and the time evolution process $e^{-i\mathcal{H}\Delta t_{\mathrm{in}}}$. For instance, if the input interval is close to zero ($\Delta t_{\mathrm{in}}\rightarrow 0$), the newly added input state is traced out almost immediately, and the distance remains largely unchanged. Conversely, a larger $\Delta t_{\mathrm{in}}$ yields more intricate dynamics, where the partial trace operation can lead to a more significant reduction in distance, the details of which depend on the Hamiltonian.
Figures S6(a) and S6(b) display the distance $||\rho_{A}^{(k)}-\rho_{B}^{(k)}||$ as a function of the input step $k$ for various input intervals $\Delta t_{\mathrm{in}}$ in the SYK4 and SYK2 models, respectively. The initial states $\rho_{A,B}^{(0)}$ are randomly sampled from the half-filling sector. To accommodate values spanning several orders of magnitude, both the mean and median distances over $500$ pairs of initial states are presented. For all models investigated, our results strongly indicate that the update rule [Eq. (2) of the main text] is contractive for random initial conditions and inputs: $||\rho_{A}^{(k+1)}-\rho_{B}^{(k+1)}||<||\rho_{A}^{(k)}-\rho_{B}^{(k)}||$. This observation suggests that ESP is satisfied for these systems—at least asymptotically as $k\rightarrow\infty$.
@@
Notably, the convergence speeds of the SYK4 and SYK2 models differ significantly. In the SYK4 model, the distance metric consistently decays exponentially toward 00, but the decay rate is sensitive to the input interval $\Delta t_{\mathrm{in}}$. For smaller intervals ($\Delta t_{\mathrm{in}}=0.1$, $0.5$), both the mean and median distances converge relatively slowly due to less effective erasure of past states upon the partial trace. Conversely, larger $\Delta t_{\mathrm{in}}$ values result in much faster convergence, with the distance behaving similarly across these larger input intervals. In stark contrast, the SYK2 model exhibits considerably slower distance decay. While its median distance also decreases exponentially, the mean distance decays only polynomially and remains relatively high. This behavior reflects the integrable nature of the SYK2 model, wherein numerous internal conserved quantities can impede convergence. These conserved quantities also generate large statistical fluctuations, which explains the disparity between the mean and median distances [Fig. S6(b)] and the large standard deviation in the performance metric (the SYK2-dominated regime in Fig. 4 of the main text).
@@
Figure S7: Median distance $||\rho_{A}^{(k)}-\rho_{B}^{(k)}||$ at $k=4000$ versus the coupling strength $\kappa_{2}/J_{4}$. Medians are calculated from $500$ independent pairs of random initial states $\rho_{A}^{(0)}$ and $\rho_{B}^{(0)}$, and are presented for two input intervals: $\Delta t_{\mathrm{in}}=1$ (purple) and $\Delta t_{\mathrm{in}}=50$ (orange). The vertical lines indicate the onset of WD (left) and Poisson (right) statistics.
@@
Figure S7 displays the median distance $||\rho_{A}^{(k)}-\rho_{B}^{(k)}||$ for the interpolated model as a function of the coupling strength $\kappa_{2}/J_{4}$. The distance is evaluated at the final step considered ($k=4000$) and is presented for two input intervals: $\Delta t_{\mathrm{in}}=1$ and $\Delta t_{\mathrm{in}}=50$. Larger final distances indicate slower convergence, whereas smaller ones suggest rapid decay. Consistent with Figs. S6(a) and S6(b), the distance approaches 00 for small $\kappa_{2}/J_{4}$ and remains relatively large in the strong $\kappa_{2}/J_{4}$ regime for both input intervals. However, the behavior differs between the two input intervals at intermediate $\kappa_{2}/J_{4}$ values. For $\Delta t_{\mathrm{in}}=1$, the distance deviates from 00 at relatively small $\kappa_{2}/J_{4}$, near the onset of WD statistics. In contrast, with $\Delta t_{\mathrm{in}}=50$, it stays close to 00 until $\kappa_{2}/J_{4}$ approaches the boundary of the Poisson regime. This comparison indicates that convergence becomes harder to achieve as the system nears the integrable regime, which requires a large $\Delta t_{\mathrm{in}}$ to generate the necessary complex dynamics.
We note that the authors in Ref. [20] discuss QRC performance in spin-based quantum reservoirs, attributing poor performance in localized phases to slow convergence. Our QRC likewise exhibits slow convergence in the SYK2-dominant regime (Fig. S7). However, a crucial difference exists: the QRC in Ref. [20] performs poorly in both linear memory and nonlinear processing within the localized phase, whereas our QRC maintains robust memory performance in the integrable regime [Fig. 3(d) of the main text]. Furthermore, for $\Delta t_{\mathrm{in}}=50$, the parametric edge of many-body quantum chaos [Fig. 4(a) of the main text] appears at $\kappa_{2}/J_{4}$ values almost an order of magnitude smaller than those at which the distance metric first deviates from 00 in Fig. S7. This disparity seems to suggest that convergence and the parametric edge are not directly related in the model under investigation.

### v3

@@ Abstract
The most pronounced trend is observed for $\Delta t_{\mathrm{in}}=50$ [Fig. 4(a)]. When $\kappa_{2}/J_{4}$ is varied, the NARMA performance exhibits a clear dip in the NMSE at a coupling strength just before the onset of the many-body quantum chaotic regime. This dip provides another manifestation of the “edge of many-body quantum chaos”, in this case in the parameter domain. By contrast, for the short input interval $\Delta t_{\mathrm{in}}=1$, varying the coupling strength yields no performance enhancement [Fig. 4(b)]. At such a short time-scale, the QRC does not reflect chaotic characteristics, so the performance changes only gradually, without any distinct peak at the chaos boundary.
Discussions. — In the strongly chaotic regime with large $\Delta t_{\mathrm{in}}$, the QRC provides a highly nonlinear mapping, but the intrinsic featurelessness of the quantum chaotic dynamics makes input-specific signatures relatively indistinct. Conversely, in the nonchaotic regime or with moderately small $\Delta t_{\mathrm{in}}$, the QRC retains input memory but sacrifices nonlinear expressivity. At the edge of many-body quantum chaos, the expressivity of the input-to-output map is optimally balanced against the preservation of input-specific features, leading to enhanced performance. From this picture, it follows that the precise location of this edge and the achievable gain depend on the task. Tasks that demand long memory shift the performance peak toward the nonchaotic side (e.g., the STM task with large $d$), whereas tasks that require stronger nonlinearity at fixed memory shift it toward the chaotic side (e.g., the nonlinear STM task). In extreme cases—when a task overwhelmingly requires either memory or nonlinearity, or when its complexity exceeds the QRC’s capacity—the edge-induced performance enhancement may not appear. See the Supplemental Material for additional performance evaluations [89].
@@
Recent advances in QRC have reported related performance enhancements, and we believe that a concise comparison highlights the significance of our contributions. One of the authors and his collaborators identified a performance peak near the operational regime boundary in their feedback-driven QRC protocol [32]; however, this enhancement appears predominantly due to classical instabilities in feedback connections, rather than intrinsically quantum phenomena. In the purely quantum regime, Martínez-Peña et al. demonstrated enhanced performance near the ergodic-localization transition in the spin-based QRC [20, 23, 33]. While this finding highlights the importance of dynamical phases in QRC, the role of many-body quantum chaos, and its relation to RMT, remain largely unexplored. Recently, Llodrà et al. analyzed QRC performance in connection with quantum chaos appearing at the quantum phase transition in a clean system, which is complementary to our study on random systems [34]. From a standpoint of temporal domain, investigations into quantum chaos remain sparse: Vetrano et al. explored the role of scrambling times, though within a different framework (quantum extreme learning machines) [94].
@@
Building upon these studies, our work directly demonstrates the relationship between quantum chaos and QRC. Using the SYK model—a canonical many-body quantum chaotic system—as a testbed, we have extended the concept of the “edge of chaos” into the quantum many-body realm by uncovering two distinct edges: one is the “temporal edge”, which emerges before the Thouless time, and the other is the “parametric edge”, which appears near the quantum chaotic-to-integrable transition. Since our RMT-based analysis makes no system-specific assumptions, we expect the observed “edge of many-body quantum chaos” to be broadly observable across platforms. Our study therefore provides clear design guidelines for engineering quantum reservoirs, unlocking advanced information processing with quantum many-body systems.
@@ I.1 Level statistics
As illustrated in Fig. 2(b) of the main text, continuously interpolating between the SYK4 and SYK2 models drives the system from a many-body quantum chaotic phase into an integrable regime. To elucidate this behavior in greater detail, Fig. S2(a) presents the full distribution $P(r)$ for several intermediate values of the coupling ratio $\kappa_{2}/J_{4}$, extending the results presented in Fig. 2(a) of the main text. As $\kappa_{2}/J_{4}$ increases, $P(r)$ evolves smoothly from the WD (GUE) statistics toward the Poisson distribution. In Fig. S2(b), we plot the mean ratio $\langle r\rangle$ as a function of $\kappa_{2}/J_{4}$ for various system sizes at $N_{p}=\lfloor N/2\rfloor$. The chaotic-integrable transition is consistently observed, although a precise determination of the boundary would require even larger system sizes [84]. For the purposes of this work, we provisionally define the boundary of the quantum chaotic regime by the coupling strength at which $\langle r\rangle$ first deviates from its RMT value. This reference point is used to discuss the parametric edge of many-body quantum chaos in Figs. 4(a) and 4(b) of the main text.
@@ I.2 Spectral form factor
Figure S3: (a), (b) The SFF for the SYK4 model and the SYK2 model, shown for various system sizes. Data are averaged over $20000$ realizations for $N=6,7,8$, $10000$ for $N=9$, and $2500$ for $N=10$. (c) The plateau time $t_{\mathrm{p}}$ of the SFF as a function of the system size NN. The blue and red dashed lines indicate reference scaling behaviors, $\mathcal{N}=\binom{N}{N_{p}}$ and $N$, respectively.
@@
where the sum runs over the eigenvalues with the particle number $N_{p}$, and $\mathcal{N}=\binom{N}{N_{p}}$ denotes the dimension of the corresponding Hilbert space. As discussed in the main text, the SFF exhibits universal characteristics in the quantum chaotic systems: at early times, it decreases due to the phase interference, while at sufficiently late times, it shows a linear ramp before the saturation to a plateau value [90, 91, 92]. In Figs. S3(a) and S3(b), we present the SFF for the SYK4 and SYK2 models, respectively, for various system sizes. The ramp-plateau behavior is clearly demonstrated especially for the former system, confirming its many-body quantum chaotic nature. In contrast, the SFF for the latter model shows the dip and plateau over a much narrower time window. Indeed, Fig. S3(c) shows that the plateau time $t_{\mathrm{p}}$ scales with the system size as $t_{\mathrm{p}}\sim\mathcal{N}$ for the SYK4 system and as $t_{\mathrm{p}}\sim N$ for the SYK2 system [73, 92]. This contrast further highlights the difference between the quantum chaotic and integrable models. We note that the Thouless time in the SYK4 model is known to scale as $t_{\mathrm{Th}}\sim\sqrt{\mathcal{N}}$ [73], but the available sizes here are insufficient to reliably extract it.
@@
Figure S4: The SFF for the rescaled Hamiltonian $\mathcal{H}\rightarrow\mathcal{H}/$$\mathcal{H}||$ at several coupling strength $\kappa_{2}$, with a fixed system size of $N=8$ and $J_{4}=1$. The dashed lines signify $t=1$ and $50$, in relation to the input interval $\Delta t_{\mathrm{in}}$ used in to evaluate QRC performance. The results are averaged over $20000$ realizations.
@@
When interpolating between the two models, the overall energy scale varies with the coupling strength. To compare QRC performance on an equal footing, we normalize the Hamiltonian by its spectral norm, $\mathcal{H}\rightarrow\mathcal{H}/$$\mathcal{H}||$. This normalization leaves the level spacing ratio invariant, as well the chaotic nature of the system. However, it introduces a uniform prefactor in the temporal domain, affecting the SFF [Eq. (S3)] and the $\Delta t_{\mathrm{in}}$ dependency of QRC performance. Figure S4 displays the SFF for the interpolated model after the rescaling. In the quantum chaotic regime ($\kappa_{2}=0.1$, 11), the typical ramp-plateau profiles emerge, while in the integrable regime ($\kappa_{2}=20$, $100$), the SFF dips and plateaus at much earlier times. Apart from the overall time-scale shift introduced by the Hamiltonian rescaling, these behaviors mirror those of the SYK4 and SYK2 models in Fig. 2(b) of the main text.
@@
In our QRC performance benchmarks in the parametric domain (Fig. 4 of the main text), we selected the input time intervals $\Delta t_{\mathrm{in}}=50$ and $\Delta t_{\mathrm{in}}=1$; the corresponding times are marked by the dashed lines in Fig. S4. For the quantum chaotic regime, the longer $\Delta t_{\mathrm{in}}$ falls within the RMT ramping window, while the shorter one precedes the dip. Similarly for the integrable models, these two times effectively sandwich the dip. These observations validate both the Hamiltonian normalization procedure and our choice of $\Delta t_{\mathrm{in}}$ for probing distinct temporal regimes relevant to QRC. It is worth noting that precisely characterizing chaos in the interpolated model via the SFF typically requires a more sophisticated procedure called unfolding [85]. Unfolding recasts the energy spectrum such that it will exhibit a constant level spacing; it is not an operation directly applied to the Hamiltonian itself. We do not employ this procedure because our primary motivation is to characterize QRC performance by focusing on the properties of the original, physically implemented Hamiltonian.
@@ II Dynamics of quantum reservoir system
Figures S5(a) and S5(b) illustrate the dynamics of read-out signals for the SYK4 and SYK2 models, respectively, for input intervals of $\Delta t_{\mathrm{in}}=1$ and $\Delta t_{\mathrm{in}}=10$. The dashed lines indicate when new inputs are provided. Markers show $\langle c_{i}^{\dagger}c_{i}\rangle$ at times $v\Delta t_{\mathrm{in}}/V$ ($v=0,\dots,V$), with the filled marker specifically denoting site $1$, i.e., the input site. The dynamics at the input site change abruptly after an input, showing a larger variation range than other sites; hence, its vertical axis is displayed separately.
With a short input interval ($\Delta t_{\mathrm{in}}=1$), each new input arrives before the previous one has fully influenced the system. Thus, the resulting dynamics is relatively simple, as shown in the upper panels of both Figs. S5(a) and S5(b). In contrast, system-dependent characteristics emerge when the input interval is longer ($\Delta t_{\mathrm{in}}=10$). In the SYK4 model [lower panel of Fig. S5(a)], all readout signals tend toward quasi-stationary values, whereas in the SYK2 model [lower panel of Fig. S5(b)], the signals exhibit continuous oscillations. Notably, although both dynamics appear sufficiently rich, their QRC performance differs significantly, as demonstrated in Figs. 3(c) and 3(d) of the main text. Because the output $y^{(k)}$ is computed from these readout signals, this performance gap must originate from the difference in dynamics. However, the underlying reason for this performance gap is not immediately obvious from the raw signals in Fig. S5.
These $\Delta t_{\mathrm{in}}$-dependent behaviors raise interest in the relationship between QRC performance and scrambling time—the timescale over which information delocalizes [76]. Since scrambling relates to the propagation of encoded input information, its connection to memory performance is naturally anticipated [35]. Accordingly, the observed $\Delta t_{\mathrm{in}}$ dependence of STM performance [upper panels of Figs. 3(c) and 3(d) of the main text] could plausibly reflect the underlying scrambling timescale. However, the pronounced enhancement in nonlinear processing performance [lower panel of Fig. 3(c) of the main text] cannot be explained by scrambling alone; scrambling itself is primarily associated with memory. Moreover, the convergence of QRC performance toward that of the Haar QRC model when $\Delta t_{\mathrm{in}}$ exceeds the Thouless time underscores a fundamental connection to RMT. These observations collectively suggest that the Thouless time plays the more decisive role in determining overall QRC performance. Nevertheless, the detailed relationship between scrambling time, Thouless time, and the temporal edge of many-body quantum chaos warrants further investigation.
@@ III Convergence property
Figure S6: Evolution of the distances $\rho_{A}^{(k)}-\rho_{B}^{(k)}||$ as a function of the step $k$ for (a) the SYK4 model and (b) the SYK2 model. The left panels show the median distance (semilog scale), and the right panels show the mean distance [semilog for (a) and log-log for (b)], calculated from $500$ independent pairs of initial states $\rho_{A}^{(0)}$ and $\rho_{B}^{(0)}$. Colors indicate the input interval $\Delta t_{\mathrm{in}}$.
[…]
@@
The convergence speeds of the SYK4 and SYK2 models differ significantly. In the SYK4 model, the distance metric consistently decays exponentially toward 0, but the decay rate is sensitive to the input interval $\Delta t_{\mathrm{in}}$. For smaller intervals ($\Delta t_{\mathrm{in}}=0.1$, $0.5$), both the mean and median distances converge relatively slowly due to less effective erasure of past states upon the partial trace. Conversely, larger $\Delta t_{\mathrm{in}}$ values result in much faster convergence, with the distance behaving similarly across these larger input intervals. In stark contrast, the SYK2 model exhibits considerably slower distance decay. While its median distance also decreases exponentially, the mean distance decays only polynomially and remains relatively high. This behavior reflects the integrable nature of the SYK2 model, wherein numerous internal conserved quantities can impede convergence. These conserved quantities also generate large statistical fluctuations, which explains the disparity between the mean and median distances [Fig. S6(b)] and the large standard deviation in the performance metric (the SYK2-dominated regime in Fig. 4 of the main text).
@@
Figure S7: Median distance $\rho_{A}^{(k)}-\rho_{B}^{(k)}||$ at $k=4000$ versus the coupling strength $\kappa_{2}/J_{4}$. Medians are calculated from $500$ independent pairs of random initial states $\rho_{A}^{(0)}$ and $\rho_{B}^{(0)}$, and are presented for two input intervals: $\Delta t_{\mathrm{in}}=1$ (purple) and $\Delta t_{\mathrm{in}}=50$ (orange). The vertical lines indicate the onset of WD (left) and Poisson (right) statistics.
@@
Figure S7 displays the median distance $\rho_{A}^{(k)}-\rho_{B}^{(k)}||$ for the interpolated model as a function of the coupling strength $\kappa_{2}/J_{4}$. The distance is evaluated at the final step considered ($k=4000$) and is presented for two input intervals: $\Delta t_{\mathrm{in}}=1$ and $\Delta t_{\mathrm{in}}=50$. Larger final distances indicate slower convergence, whereas smaller ones suggest rapid decay. Consistent with Figs. S6(a) and S6(b), the distance approaches 0 for small $\kappa_{2}/J_{4}$ and remains relatively large in the strong $\kappa_{2}/J_{4}$ regime for both input intervals. However, the behavior differs between the two input intervals at intermediate $\kappa_{2}/J_{4}$ values. For $\Delta t_{\mathrm{in}}=1$, the distance deviates from 0 at relatively small $\kappa_{2}/J_{4}$, near the onset of WD statistics. In contrast, with $\Delta t_{\mathrm{in}}=50$, it stays close to 0 until $\kappa_{2}/J_{4}$ approaches the boundary of the Poisson regime. This comparison indicates that convergence becomes harder to achieve as the system nears the integrable regime, which requires a large $\Delta t_{\mathrm{in}}$ to generate the necessary complex dynamics.
We note that Ref. [20] analyzes QRC performance in spin-based quantum reservoirs, attributing the poor performance observed in localized phases to slow convergence. Our QRC likewise exhibits slow convergence in the SYK2-dominant regime (Fig. S7). Nevertheless, as shown in Fig. 4(a) of the main text, the parametric edge of many-body quantum chaos appears at $\kappa_{2}/J_{4}$ values almost an order of magnitude smaller than those at which the distance metric first deviates from 0 in Fig. S7. Moreover, although the convergence metric varies significantly with $\kappa_{2}/J_{4}$ even for $\Delta t_{\mathrm{in}}=1$, the performance peak associated with the parametric edge does not appear in Fig. 4(b) of the main text. Taken together, these observations suggest that convergence and the parametric edge are not directly related here.